# Propensity Score Methods for Local Test Score Equating: Stratification and Inverse Probability Weighting

### v1

@@ 6.3 Results
Figure 7 shows the bias and RMSE under the weak correlation scenario, for selected $\theta$ values across the score range, comparing the performance of the three considered equating methods. What is immediately noticeable is that the bias and RMSE are not displayed for all score values. The reason is that all results for score values with a probability lower than $10^{-4}$ have been omitted as their equatings are not considered stable enough. In practice, this choice does not have any consequences since only very few examinees would obtain one of the omitted values. These findings aligns well with what was found in Wiberg and van der Linden, (2011), who pointed out that these scores should never be reported without a warning about their lack of accuracy.
In Figure 7 panel (a), the anchor method demonstrates relatively stable bias across the score range, with a decreasing bias for higher values of $\theta$. A similar trend is seen for the IPW method, but with a consistently smaller bias compared to the anchor method. For the propensity score stratification method, the bias varies more across the score range but with a similar shape for all values of $\theta$. Panel (b) in Figure 7 displays the RMSE. The RMSEs are very similar to the bias values, indicating only very small variability of the estimated equated score across simulations. One clear difference is that the RMSE of the anchor method is even more similar to that of the IPW method.
(a) The bias.
(b) The RMSE.
@@
Figure 7: The performance in terms of average bias and RMSE of anchor-based, propensity score-based, and IPW-based equating methods across sum scores for selected theta values for $N=1000$ and $J=40$ when the correlations between the covariates and the test scores are of low strength.
In Figure 8, panel (a) shows the bias for the anchor, propensity score stratification, and IPW equating methods under the medium correlation scenario. Similarly to the weak correlation scenario, the bias for the anchor method is quite stable across the score range for a given value of $\theta$. As $\theta$ increases, the bias steadily decreases for the anchor method. This is true for the propensity score stratification method as well, although the bias and RMSE (panel (b)) is increasing across the score scale fora given value of $\theta$. In contrast, the bias and RMSE of the IPW method is rather stable for all values of $\theta$. We also conclude that its performance is, surprisingly, slightly worse when the covariates have a stronger correlation to the test scores.
(a) The bias.
(b) The RMSE.
@@ 7 Discussion
From the empirical analysis using data from the SweSAT, we observed that both propensity score-based methods produced equated scores that varied depending on the percentile of the propensity score conditioned upon. This result is in line with the results of previous local equating studies when other proxies have been used to capture differences between group differences (e.g. van der Linden and Wiberg, , 2010; Wiberg and van der Linden, , 2011; Wiberg et al., , 2014). This variation was particularly noticeable at the lower and higher ends of the score scale, while the equated scores were relatively similar in the mid-range. The anchor-based method exhibited a similar pattern, suggesting that conditioning on either an anchor score or estimated propensity scores can capture variations in examinee abilities across different score levels. The propensity score stratification method showed equated scores that were closely aligned for the 30th, 50th, and 90th percentiles, indicating consistency in the mid to higher ability ranges. The IPW method also demonstrated stability in the mid-range but showed differences at the extremes of the score distribution.
In the simulation study, we evaluated the performance of the proposed methods under various conditions, varying factors such as the strength of the correlation between covariates and the latent ability, and the sample size. Note that, to examine different levels of correlation has neither been examined in local equating studies (e.g. van der Linden and Wiberg, , 2010; Wiberg and van der Linden, , 2011), nor when equating with propensity scores (Wallin and Wiberg, , 2019). The results indicated that when the correlation between covariates and the test scores was weak, the IPW method consistently exhibited smaller bias compared to the anchor-based and propensity score stratification methods across most of the score range. The bias and RMSE for the IPW method were relatively stable, suggesting robustness in scenarios where covariates are not strongly predictive of the latent ability. When the correlation between covariates and test scores was medium, the performance of the anchor-based method improved, showing decreased bias at higher levels of latent ability. The propensity score stratification method exhibited increasing bias and RMSE across the score scale for a given ability level, but performed better than the IPW method. Surprisingly, the IPW method’s performance was slightly worse in the medium correlation scenario compared to the weak correlation scenario, although it still maintained relatively low bias and RMSE.
@@
The results align with previous research that has explored the use of covariates and propensity scores in equating. Wiberg and Bränberg, (2015) demonstrated that covariate-adjusted equating methods could effectively account for group differences in the NEC design. Our study extends this work by applying propensity score methods within the local equating framework, addressing the challenge of fulfilling Lord’s equity requirement in the absence of an anchor test. Additionally, our findings corroborate those of Wallin and Wiberg, (2019), who proposed the use of propensity scores in the NEC design and highlighted the potential of these methods to reduce equating bias.
Despite the promising results, several limitations should be acknowledged. First, the success of propensity score methods relies heavily on the quality and relevance of the covariates used. In practice, covariates must be related to the latent ability to effectively adjust for group differences. If important covariates are omitted or the relationship between covariates and ability is too weak, the equating transformation may be biased. Additionally, the assumption of unconfoundedness - that all relevant covariates have been included and correctly measured - is a strong one and may not hold in all testing scenarios. Thus it is important to perform robustness studies, similar to Wallin and Wiberg, (2023).

### v2

@@ 6.3 Results
Figure 8 shows the bias and RMSE under the weak correlation scenario, for selected $\theta$ values across the score range, comparing the performance of the three considered equating methods. What is immediately noticeable is that the bias and RMSE are not displayed for all score values. The reason is that all results for score values with a probability lower than $10^{-4}$ have been omitted as their equatings are not considered stable enough. In practice, this choice does not have any consequences since only very few examinees would obtain one of the omitted values. These findings align well with what was found in Wiberg and van der Linden (2011), who pointed out that these scores should never be reported without a warning about their lack of accuracy. As noted by these authors, such scores fall in regions where the equating transformation becomes highly unreliable, making unconditional reporting problematic. In practice, this choice affects very few test takers since only a very small number would obtain one of these extreme values, but we argue that the potential for substantial equating error in these regions justifies their exclusion from routine reporting.
In Figure 8, the three methods yield very similar results across most of the score range, especially the propensity score stratification and IPW methods, which are nearly indistinguishable in both bias and RMSE. The average absolute difference between these two methods is only 0.08 for both metrics, with no substantial divergence at any particular score level. The anchor method also performs similarly in many regions. Differences emerge for certain scores, especially at the lower end of the score range where the anchor method exhibits higher bias and RMSE. Unlike the propensity score-based methods, which directly account for covariate information, the anchor method relies solely on the assumption that examinees with the same anchor scores have similar abilities across forms. When the correlation between covariates and ability weakens, the conditional distribution of abilities given anchor scores may differ more substantially across the two populations, particularly at the extremes of the score distribution. This explains why the anchor method’s performance is slightly worse at the lower end of the score range, where sample sizes are typically smaller and estimation is inherently more challenging. Interestingly, despite not directly incorporating covariates in its equating procedure, the anchor method’s performance is indirectly affected by covariate correlation changes through the form assignment mechanism. When covariates have weaker correlations with ability, the propensity model for form assignment becomes more heavily influenced by the anchor score alone, potentially creating more systematic differences in ability distributions between forms that are inadequately captured by the anchor items. This highlights an important characteristic of non-equivalent group equating designs: the anchor method’s effectiveness depends not only on the anchor items themselves but also on the underlying mechanisms determining sample selection into different forms.
@@
Figure 8: The performance in terms of average bias and RMSE of anchor-based, propensity score-based, and IPW-based equating methods across sum scores for selected theta values for $N=1000$ and $J=40$ when the correlations between the covariates and the test scores are of weak strength.
In Figure 9, panel (a) shows the bias for the anchor, propensity score stratification, and IPW equating methods under the medium correlation scenario. The results are similar to the weak correlation setting, however, both bias and RMSE are slightly smaller which is to be expected.
@@ 7 Discussion
From the empirical analysis using data from the SweSAT, we observed that both propensity score-based methods produced equated scores that varied depending on the percentile of the propensity score conditioned upon. This result is in line with the results of previous local equating studies when other proxies have been used to capture differences between group differences (e.g. van der Linden and Wiberg, 2010; Wiberg and van der Linden, 2011; Wiberg et al., 2014). This variation was particularly noticeable at the lower and higher ends of the score scale, while the equated scores were relatively similar in the mid-range. The anchor-based method exhibited a similar pattern, suggesting that conditioning on either an anchor score or estimated propensity scores can capture variations in examinee abilities across different score levels. The propensity score stratification method showed equated scores that were closely aligned for the 30th, 50th, and 90th percentiles, indicating consistency in the mid to higher ability ranges. The IPW method also demonstrated stability in the mid-range but showed differences at the extremes of the score distribution.
In the simulation study, we evaluated the performance of the proposed methods under various conditions, varying factors such as the strength of the correlation between covariates and the latent ability, and the sample size. Note that, to examine different levels of correlation has neither been examined in local equating studies (e.g. van der Linden and Wiberg, 2010; Wiberg and van der Linden, 2011), nor when equating with propensity scores (Wallin and Wiberg, 2019). The results indicated that when the correlation between covariates and the test scores was weak, the IPW method consistently exhibited smaller bias compared to the anchor-based and propensity score stratification methods across most of the score range. The bias and RMSE for the IPW method were relatively stable, suggesting robustness in scenarios where covariates are not strongly predictive of the latent ability. When the correlation between covariates and test scores was medium, the performance of the anchor-based method improved, showing decreased bias at higher levels of latent ability. The propensity score stratification method exhibited increasing bias and RMSE across the score scale for a given ability level, but performed better than the IPW method. Surprisingly, the IPW method’s performance was slightly worse in the medium correlation scenario compared to the weak correlation scenario, although it still maintained relatively low bias and RMSE.
@@
The results align with previous research that has explored the use of covariates and propensity scores in equating. Wiberg and Bränberg (2015) demonstrated that covariate-adjusted equating methods could effectively account for group differences in the NEC design. Our study extends this work by applying propensity score methods within the local equating framework, addressing the challenge of fulfilling Lord’s equity requirement in the absence of an anchor test. Additionally, our findings corroborate those of Wallin and Wiberg (2019), who proposed the use of propensity scores in the NEC design and highlighted the potential of these methods to reduce equating bias.
Although local equating procedures may produce different conversion functions for different subgroups the idea of stratified reporting is not new in testing programs as several of them already report scores conditional on subgroup membership (e.g., by grade level, test version, or test-taking language). Also, our proposed local equating methods can be implemented to obtain conversion functions for a manageable number of strata, rather than a separate function for every individual. The conversion tables can be pre-computed and embedded in scoring software, making operational implementation feasible. Finally, as local equating can reduce bias and improve score interpretation in the presence of group differences this can outweigh the added complexity of using local equating.
Despite the promising results in this study, several limitations should be acknowledged. First, the success of propensity score methods relies heavily on the quality and relevance of the covariates used. In practice, covariates must be related to the latent ability to effectively adjust for group differences. If important covariates are omitted or the relationship between covariates and ability is too weak, the equating transformation may be biased. Additionally, the assumption of unconfoundedness - that all relevant covariates have been included and correctly measured - is a strong one and may not hold in all testing scenarios. Thus it is important to perform robustness studies, similar to Wallin and Wiberg (2023).